OpenStudy (anonymous):

Solve. 2x + 6 – 6x = 30

OpenStudy (anonymous):

2x-6x=30-6 -4x=24 x=24/-4 x=-6

OpenStudy (anonymous):

Solve. 4(x – 3) + 5x = –48

OpenStudy (anonymous):

But give me a medal

OpenStudy (anonymous):

4x-12+5x=-48 4x+5x=-48+12 9x=-36 x=-36/9 x=-4

OpenStudy (anonymous):

medall!!!

OpenStudy (anonymous):

there can u get a couple more for me

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Frank’s quarter grades are 82.5, 94.7, and 87.9. What grade, rounded to the nearest percent, will Frank have to earn in the last quarter to receive at least a 90% final course grade?

OpenStudy (anonymous):

open study is suposed to help people not giving answer i will give you the equation that you must use to find the answer its so easy 

OpenStudy (anonymous):